OpenStudy (anonymous):

Find all solutions to the equation sin^2(x)cos^2(x) = (2 - √2) / 16

OpenStudy (experimentx):

change sin in to cos or change cos into sin, and simplify you will have a quadratic equation in sin square or cos square solve it, and then find x

OpenStudy (anonymous):

i think that might work, but you get a polynomial of degree 4 which will be fairly annoying to solve

OpenStudy (anonymous):

maybe try \[\frac{1}{4}\sin(2x)=\frac{2+\sqrt{2}}{16}\] and go from there. not sure which is easiest

OpenStudy (mertsj):

No doubt. Thank goodness you've arrived.

OpenStudy (anonymous):

well that was wrong. i mean \[\frac{1}{4}\sin^2(2x)=\frac{2+\sqrt{2}}{16}\]

OpenStudy (anonymous):

still a pain. there is probably some identity that i am missing that would work better. but at least this one gives you \[\sin^2(2x)=\frac{2+\sqrt{2}}{4}\]

OpenStudy (anonymous):

off by a sign it is \[\sin^2(2x)=\frac{2-\sqrt{2}}{4}\]

OpenStudy (anonymous):

I got one answer as pi/16 and I am not sure if that is correct

OpenStudy (anonymous):

we could check it i guess.

OpenStudy (anonymous):

dang you are right. how did you get it?

OpenStudy (anonymous):

i am betting experimentx method is actually easiest

OpenStudy (experimentx):

i think your method is a lot simpler and better than mine.

OpenStudy (anonymous):

well it is still a pain either way

OpenStudy (experimentx):

a special case ..!

OpenStudy (anonymous):

I got pi/16 because in the beginning I just sqrt both sides and I was left with sinxcosx=sqrt(2-sqrt(2)/4 and then I just went from there

OpenStudy (anonymous):

ok but don't forget \[\sin(x)\cos(x)=-\frac{2-\sqrt{2}}{16}\]

OpenStudy (mertsj):

\[(\frac{1-\cos(2x)}{2})(\frac{1+\cos(2x)}{2})=\frac{2=\sqrt{2}}{16}\]

OpenStudy (mertsj):

\[\frac{1-\cos^2(2x)}{4}=\frac{2-\sqrt{2}}{16}\]

OpenStudy (mertsj):

\[\sin^2(2x)=\frac{2-\sqrt{2}}{4}\]

OpenStudy (mertsj):

But \[\sin^2\frac{u}{2}=\frac{1-cosu}{2}\]

OpenStudy (mertsj):

So \[\sin ^2(\frac{4x}{2})=\frac{1-\cos (4x)}{2}\]= \[\frac{2-\sqrt{2}}{4}=\frac{1}{2}-\frac{\frac{\sqrt{2}}{2}}{2}\] So \[\cos (4x)=\frac{\sqrt{2}}{2}\] \[4x=\frac{\pi}{4},\frac{7\pi}{4},\frac{9\pi}{4},\frac{15\pi}{4}, \frac{17\pi}{4},\frac{23\pi}{4},\frac{25\pi}{4}, \frac{31\pi}{4}\]

OpenStudy (mertsj):

\[x=\frac{\pi}{16},\frac{7\pi}{16},\frac{9\pi}{16},\frac{15\pi}{16},\frac{17\pi}{16},\frac{23\pi}{16},\frac{25\pi}{16},\frac{31\pi}{16} +2n \pi\]